Each of the two congruent sides of an isosceles triangle is 4 cm longer than half the base. Find the length of each side of the triangle if its perimeter is 36 cm.
Answer by stanbon(75887)

Each of the two congruent sides of an isosceles triangle is 4 cm longer than half the base. Find the length of each side of the triangle if its perimeter
is 36 cm.
-----
Let the base be "x":
Each of the equal sides must be (x/2)+4:
-----
Equation:
x + (x/2)+4 + (x/2)+4 = 36
2x + 8 = 36
x = 14 cm (base)
-------
x/2 + 4 = 11 cm (each of the equal sides)
==========
Cheers,
Stan H.
-----